# Blameless Users in a Clean Room: Defining Copyright Protection for Generative Models

### v1

@@ B.4 On NAF’s legal relevance
With the above in mind, [VKB23] explain the relevance of NAF to copyright liability.
@@
As for why it’s a stronger guarantee, the argument is as follows. The probability that the defendant’s work—produced by the real model $p$—is substantially similar to plaintiff’s work is not much greater than the probability would have been had the defendant used the safe model (which had no access). We heuristically expect the latter probability to be miniscule. Hence, substantial similarity between the defendant’s and plaintiff’s works is exceedingly unlikely.
Elkin-Koren et al. [EKHLM24] correctly argue that copyright cannot be “reduced to privacy.” However, this is a straw man of version of [VKB23] and the present paper (see Appendix A for additional discussion).
The sharpest criticism is by Lee, Cooper, and Grimmelman who argue that NAF is simply wrong on the law [LCG24]. “It is not a defense to copyright infringement that you would have copied the work from somewhere else if you hadn’t copied it from the plaintiff.”
@@
Appendix C Differential privacy background [DMNS06]
@@
Let $\mathsf{id}=\mathsf{ideas}(w_{1,1})$. Define $\tilde{\mathsf{u}}=\mathsf{u}_{\mathsf{id}}$ to be the user with $\mathsf{id}$ hard-coded that simulates $\mathsf{u}(\mathsf{id})$. By construction, $\tilde{\tau}(E;\mathsf{aux})=\tau(E;\mathsf{id})$ for all $\mathsf{aux}$ and all events $E$. Because $\tilde{\mathsf{u}}$ ignores its auxiliary input, we can invoke Lemma D.2. Thus, there exists $\tilde{D}$ such that $\tilde{\mathsf{u}}$ is $\beta$-blameless with respect to $\tilde{D}$, $\tilde{\mathcal{C}}=\tilde{D}$, and $\mathsf{Train}$ (condition (i) above). Moreover, $w_{i,1}\in\tilde{D}$ for some $i\in[n]$.

### v2

@@ B.4 On NAF’s legal relevance
With the above in mind, [VyasKB23] explain the relevance of NAF to copyright liability.
@@
As for why it’s a stronger guarantee, the argument is as follows. The probability that the defendant’s work—produced by the real model pp—is substantially similar to plaintiff’s work is not much greater than the probability would have been had the defendant used the safe model (which had no access). We heuristically expect the latter probability to be miniscule. Hence, substantial similarity between the defendant’s and plaintiff’s works is exceedingly unlikely.
Elkin-Koren et al. [elkinkoren2024copyright] correctly argue that copyright cannot be “reduced to privacy.” However, this is a straw man of version of [VyasKB23] and the present paper (see Appendix A for additional discussion).
The sharpest criticism is by Lee, Cooper, and Grimmelman who argue that NAF is simply wrong on the law [lee2024talkin]. “[NAF] is explicitly inspired by copyright’s concept of access, but copyright law itself does not work that way. Just as two authors can independently create identical works and each hold a copyright in theirs, it is not a defense to copyright infringement that you would have copied the work from somewhere else if you hadn’t copied it from the plaintiff.”
@@
Appendix C Differential privacy background [dwork2006calibrating]
@@
Let $\mathsf{id}=\mathsf{ideas}(w_{1,1})$. Define $\tilde{\mathsf{u}}=\mathsf{u}_{\mathsf{id}}$ to be the user with $\mathsf{id}$ hard-coded that simulates $\mathsf{u}(\mathsf{id})$, always ignoring its auxiliary input $\mathsf{aux}$. By construction, $\tilde{\tau}(E;\mathsf{aux})=\tau(E;\mathsf{id})$ for all $\mathsf{aux}$ and all events $E$. That is, the event $E$ is independent of $\mathsf{aux}$. Because $\tilde{\mathsf{u}}$ ignores its auxiliary input, we can invoke Lemma D.2. Thus, there exists $\tilde{D}$ such that $\tilde{\mathsf{u}}$ is $\beta$-blameless with respect to $\tilde{D}$, $\tilde{\mathcal{C}}=\tilde{D}$, and $\mathsf{Train}$ (condition (i) above). Moreover, $w_{i,1}\in\tilde{D}$ for some $i\in[n]$.